\begin{thebibliography}{45}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Afessa et~al.(2005)Afessa, Keegan, Gajic, Hubmayr, and
  Peters]{afessa2005influence}
Bekele Afessa, Mark~T Keegan, Ognjen Gajic, Rolf~D Hubmayr, and Steve~G Peters.
\newblock The influence of missing components of the acute physiology score of
  apache iii on the measurement of icu performance.
\newblock \emph{Intensive care medicine}, 31\penalty0 (11):\penalty0
  1537--1543, 2005.

\bibitem[Agniel et~al.(2018)Agniel, Kohane, and Weber]{agniel2018biases}
Denis Agniel, Isaac~S Kohane, and Griffin~M Weber.
\newblock Biases in electronic health record data due to processes within the
  healthcare system: retrospective observational study.
\newblock \emph{Bmj}, 361:\penalty0 k1479, 2018.

\bibitem[Agor et~al.(2019)Agor, {\"O}zalt{\i}n, Ivy, Capan, Arnold, and
  Romero]{agor2019value}
Joseph Agor, Osman~Y {\"O}zalt{\i}n, Julie~S Ivy, Muge Capan, Ryan Arnold, and
  Santiago Romero.
\newblock The value of missing information in severity of illness score
  development.
\newblock \emph{Journal of biomedical informatics}, 97:\penalty0 103255, 2019.

\bibitem[Baglin et~al.(2006)Baglin, Barrowcliffe, Cohen, Greaves, and for
  Standards~in Haematology]{baglin2006guidelines}
T~Baglin, TW~Barrowcliffe, A~Cohen, Michael Greaves, and British~Committee for
  Standards~in Haematology.
\newblock Guidelines on the use and monitoring of heparin.
\newblock \emph{British journal of haematology}, 133\penalty0 (1):\penalty0
  19--34, 2006.

\bibitem[Beaulieu-Jones et~al.(2018)Beaulieu-Jones, Lavage, Snyder, Moore,
  Pendergrass, and Bauer]{beaulieu2018characterizing}
Brett~K Beaulieu-Jones, Daniel~R Lavage, John~W Snyder, Jason~H Moore, Sarah~A
  Pendergrass, and Christopher~R Bauer.
\newblock Characterizing and managing missing structured data in electronic
  health records: data analysis.
\newblock \emph{JMIR medical informatics}, 6\penalty0 (1):\penalty0 e11, 2018.

\bibitem[Brigden(1999)]{brigden1999clinical}
Malcolm~L Brigden.
\newblock Clinical utility of the erythrocyte sedimentation rate.
\newblock \emph{American family physician}, 60\penalty0 (5):\penalty0
  1443--1450, 1999.

\bibitem[Buck(1960)]{buck1960method}
Samuel~F Buck.
\newblock A method of estimation of missing values in multivariate data
  suitable for use with an electronic computer.
\newblock \emph{Journal of the Royal Statistical Society: Series B
  (Methodological)}, 22\penalty0 (2):\penalty0 302--306, 1960.

\bibitem[Buuren and Groothuis-Oudshoorn(2010)]{buuren2010mice}
S~van Buuren and Karin Groothuis-Oudshoorn.
\newblock mice: Multivariate imputation by chained equations in r.
\newblock \emph{Journal of statistical software}, pages 1--68, 2010.

\bibitem[Che et~al.(2018)Che, Purushotham, Cho, Sontag, and
  Liu]{che2018recurrent}
Zhengping Che, Sanjay Purushotham, Kyunghyun Cho, David Sontag, and Yan Liu.
\newblock Recurrent neural networks for multivariate time series with missing
  values.
\newblock \emph{Scientific reports}, 8\penalty0 (1):\penalty0 6085, 2018.

\bibitem[Data(2016)]{data2016secondary}
MIT~Critical Data.
\newblock \emph{Secondary Analysis of Electronic Health Records}.
\newblock Springer International Publishing, 2016.

\bibitem[Duff et~al.(2018)Duff, Solis-Trapala, Driskell, Holland, Wright,
  Waldron, Ford, Scargill, Tran, Hanna, et~al.]{duff2018frequency}
Christopher~J Duff, Ivonne Solis-Trapala, Owen~J Driskell, David Holland, Helen
  Wright, Jenna~L Waldron, Clare Ford, Jonathan~J Scargill, Martin Tran,
  Fahmy~WF Hanna, et~al.
\newblock The frequency of testing for glycated haemoglobin, hba1c, is linked
  to the probability of achieving target levels in patients with suboptimally
  controlled diabetes mellitus.
\newblock \emph{Clinical Chemistry and Laboratory Medicine (CCLM)}, 57\penalty0
  (2):\penalty0 296--304, 2018.

\bibitem[Ernst et~al.(2005)Ernst, Geurts, and Wehenkel]{ernst2005tree}
Damien Ernst, Pierre Geurts, and Louis Wehenkel.
\newblock Tree-based batch mode reinforcement learning.
\newblock \emph{Journal of Machine Learning Research}, 6\penalty0
  (Apr):\penalty0 503--556, 2005.

\bibitem[Fincher and Page(1986)]{fincher1986clinical}
Ruth-Marie~E Fincher and Malcolm~I Page.
\newblock Clinical significance of extreme elevation of the erythrocyte
  sedimentation rate.
\newblock \emph{Archives of internal medicine}, 146\penalty0 (8):\penalty0
  1581--1583, 1986.

\bibitem[Fushiki(2011)]{fushiki2011estimation}
Tadayoshi Fushiki.
\newblock Estimation of prediction error by using k-fold cross-validation.
\newblock \emph{Statistics and Computing}, 21\penalty0 (2):\penalty0 137--146,
  2011.

\bibitem[Gelman and Hill(2006)]{gelman2006data}
Andrew Gelman and Jennifer Hill.
\newblock \emph{Data analysis using regression and multilevel/hierarchical
  models}.
\newblock Cambridge university press, 2006.

\bibitem[Gottesman et~al.(2018)Gottesman, Johansson, Meier, Dent, Lee,
  Srinivasan, Zhang, Ding, Wihl, Peng, et~al.]{gottesman2018evaluating}
Omer Gottesman, Fredrik Johansson, Joshua Meier, Jack Dent, Donghun Lee,
  Srivatsan Srinivasan, Linying Zhang, Yi~Ding, David Wihl, Xuefeng Peng,
  et~al.
\newblock Evaluating reinforcement learning algorithms in observational health
  settings.
\newblock \emph{arXiv preprint arXiv:1805.12298}, 2018.

\bibitem[Gottesman et~al.(2019)Gottesman, Johansson, Komorowski, Faisal,
  Sontag, Doshi-Velez, and Celi]{gottesman2019guidelines}
Omer Gottesman, Fredrik Johansson, Matthieu Komorowski, Aldo Faisal, David
  Sontag, Finale Doshi-Velez, and Leo~Anthony Celi.
\newblock Guidelines for reinforcement learning in healthcare.
\newblock \emph{Nat Med}, 25\penalty0 (1):\penalty0 16--18, 2019.

\bibitem[Ha and Regal(2016)]{ha2016anticoagulation}
Nghi~B Ha and Randolph~E Regal.
\newblock Anticoagulation in patients with cirrhosis: caught between a
  rock-liver and a hard place.
\newblock \emph{Annals of Pharmacotherapy}, 50\penalty0 (5):\penalty0 402--409,
  2016.

\bibitem[Hastie et~al.(2005)Hastie, Tibshirani, Friedman, and
  Franklin]{hastie2005elements}
Trevor Hastie, Robert Tibshirani, Jerome Friedman, and James Franklin.
\newblock The elements of statistical learning: data mining, inference and
  prediction.
\newblock \emph{The Mathematical Intelligencer}, 27\penalty0 (2):\penalty0
  83--85, 2005.

\bibitem[Hylek et~al.(2003)Hylek, Regan, Henault, Gardner, Chan, Singer, and
  Barry]{hylek2003challenges}
Elaine~M Hylek, Susan Regan, Lori~E Henault, Margaret Gardner, Andrew~T Chan,
  Daniel~E Singer, and Michael~J Barry.
\newblock Challenges to the effective use of unfractionated heparin in the
  hospitalized management of acute thrombosis.
\newblock \emph{Archives of internal medicine}, 163\penalty0 (5):\penalty0
  621--627, 2003.

\bibitem[Jagannatha et~al.(2018)Jagannatha, Thomas, and Yu]{Jagannatha2018}
Abhyuday Jagannatha, Philip Thomas, and Hong Yu.
\newblock Towards high confidence off-policy reinforcement learning for
  clinical applications.
\newblock \emph{CausalML workshop, ICML 2018.}, 2018.
\newblock URL \url{http://www-all.cs.umass.edu/pubs/2018/Jagannatha et al -
  Towards High Confidence.pdf}.

\bibitem[Jeter et~al.(2019)Jeter, Josef, Shashikumar, and
  Nemati]{jeter2019does}
Russell Jeter, Christopher Josef, Supreeth Shashikumar, and Shamim Nemati.
\newblock Does the" artificial intelligence clinician" learn optimal treatment
  strategies for sepsis in intensive care?
\newblock \emph{arXiv preprint arXiv:1902.03271}, 2019.

\bibitem[Jiang and Li(2015)]{jiang2015doubly}
Nan Jiang and Lihong Li.
\newblock Doubly robust off-policy value evaluation for reinforcement learning.
\newblock \emph{arXiv preprint arXiv:1511.03722}, 2015.

\bibitem[Komorowski et~al.(2018)Komorowski, Celi, Badawi, Gordon, and
  Faisal]{komorowski2018artificial}
Matthieu Komorowski, Leo~A Celi, Omar Badawi, Anthony~C Gordon, and A~Aldo
  Faisal.
\newblock The artificial intelligence clinician learns optimal treatment
  strategies for sepsis in intensive care.
\newblock \emph{Nature Medicine}, 24\penalty0 (11):\penalty0 1716, 2018.

\bibitem[Lipton et~al.(2016)Lipton, Kale, and Wetzel]{lipton2016directly}
Zachary~C Lipton, David Kale, and Randall Wetzel.
\newblock Directly modeling missing data in sequences with rnns: Improved
  classification of clinical time series.
\newblock In \emph{Machine Learning for Healthcare Conference}, pages 253--270,
  2016.

\bibitem[Lopez-Martinez et~al.(2019)Lopez-Martinez, Eschenfeldt, Ostvar,
  Ingram, Hur, and Picard]{lopez2019deep}
Daniel Lopez-Martinez, Patrick Eschenfeldt, Sassan Ostvar, Myles Ingram, Chin
  Hur, and Rosalind Picard.
\newblock Deep reinforcement learning for optimal critical care pain management
  with morphine using dueling double-deep q networks.
\newblock \emph{arXiv preprint arXiv:1904.11115}, 2019.

\bibitem[Lowe et~al.(2009)Lowe, Ferris, Hernandez, and Weber]{lowe2009stride}
Henry~J Lowe, Todd~A Ferris, Penni~M Hernandez, and Susan~C Weber.
\newblock Stride--an integrated standards-based translational research
  informatics platform.
\newblock In \emph{AMIA Annual Symposium Proceedings}, volume 2009, page 391.
  American Medical Informatics Association, 2009.

\bibitem[Mann and Whitney(1947)]{mann1947test}
Henry~B Mann and Donald~R Whitney.
\newblock On a test of whether one of two random variables is stochastically
  larger than the other.
\newblock \emph{The annals of mathematical statistics}, pages 50--60, 1947.

\bibitem[Miettinen(1974)]{miettinen1974confounding}
Olli Miettinen.
\newblock Confounding and effect-modification.
\newblock \emph{American Journal of Epidemiology}, 100\penalty0 (5):\penalty0
  350--353, 1974.

\bibitem[Nemati et~al.(2016)Nemati, Ghassemi, and Clifford]{nemati2016optimal}
Shamim Nemati, Mohammad~M Ghassemi, and Gari~D Clifford.
\newblock Optimal medication dosing from suboptimal clinical examples: A deep
  reinforcement learning approach.
\newblock In \emph{2016 38th Annual International Conference of the IEEE
  Engineering in Medicine and Biology Society (EMBC)}, pages 2978--2981. IEEE,
  2016.

\bibitem[Newsome et~al.(2018)Newsome, Cramb, Davison, Dillon, Foulerton,
  Godfrey, Hall, Harrower, Hudson, Langford, et~al.]{newsome2018guidelines}
Philip~N Newsome, Rob Cramb, Suzanne~M Davison, John~F Dillon, Mark Foulerton,
  Edmund~M Godfrey, Richard Hall, Ulrike Harrower, Mark Hudson, Andrew
  Langford, et~al.
\newblock Guidelines on the management of abnormal liver blood tests.
\newblock \emph{Gut}, 67\penalty0 (1):\penalty0 6--19, 2018.

\bibitem[Pedregosa et~al.(2011)Pedregosa, Varoquaux, Gramfort, Michel, Thirion,
  Grisel, Blondel, Prettenhofer, Weiss, Dubourg, Vanderplas, Passos,
  Cournapeau, Brucher, Perrot, and Duchesnay]{scikit-learn}
F.~Pedregosa, G.~Varoquaux, A.~Gramfort, V.~Michel, B.~Thirion, O.~Grisel,
  M.~Blondel, P.~Prettenhofer, R.~Weiss, V.~Dubourg, J.~Vanderplas, A.~Passos,
  D.~Cournapeau, M.~Brucher, M.~Perrot, and E.~Duchesnay.
\newblock Scikit-learn: Machine learning in {P}ython.
\newblock \emph{Journal of Machine Learning Research}, 12:\penalty0 2825--2830,
  2011.

\bibitem[Peng et~al.(2018)Peng, Ding, Wihl, Gottesman, Komorowski, Lehman,
  Ross, Faisal, and Doshi-Velez]{peng2018improving}
Xuefeng Peng, Yi~Ding, David Wihl, Omer Gottesman, Matthieu Komorowski,
  Li-wei~H Lehman, Andrew Ross, Aldo Faisal, and Finale Doshi-Velez.
\newblock Improving sepsis treatment strategies by combining deep and
  kernel-based reinforcement learning.
\newblock In \emph{AMIA Annual Symposium Proceedings}, volume 2018, page 887.
  American Medical Informatics Association, 2018.

\bibitem[Pivovarov et~al.(2014)Pivovarov, Albers, Sepulveda, and
  Elhadad]{pivovarov2014identifying}
Rimma Pivovarov, David~J Albers, Jorge~L Sepulveda, and No{\'e}mie Elhadad.
\newblock Identifying and mitigating biases in ehr laboratory tests.
\newblock \emph{Journal of biomedical informatics}, 51:\penalty0 24--34, 2014.

\bibitem[Prasad et~al.(2017)Prasad, Cheng, Chivers, Draugelis, and
  Engelhardt]{prasad2017reinforcement}
Niranjani Prasad, Li-Fang Cheng, Corey Chivers, Michael Draugelis, and
  Barbara~E Engelhardt.
\newblock A reinforcement learning approach to weaning of mechanical
  ventilation in intensive care units.
\newblock \emph{arXiv preprint arXiv:1704.06300}, 2017.

\bibitem[Price et~al.(2013)Price, Jin, Nguyen, Krishnan, Bowen, and
  Zehnder]{price2013discordant}
Elizabeth~A Price, Jing Jin, Huong~Marie Nguyen, Gomathi Krishnan, Raffick
  Bowen, and James~L Zehnder.
\newblock Discordant aptt and anti-xa values and outcomes in hospitalized
  patients treated with intravenous unfractionated heparin.
\newblock \emph{Annals of Pharmacotherapy}, 47\penalty0 (2):\penalty0 151--158,
  2013.

\bibitem[Roderick et~al.(2002)Roderick, Little, and
  Rubin]{roderick2002statistical}
J~Roderick, A~Little, and Donald~B Rubin.
\newblock \emph{Statistical analysis with missing data}.
\newblock J. Wiley, 2002.

\bibitem[Rubin(1976)]{rubin1976inference}
Donald~B Rubin.
\newblock Inference and missing data.
\newblock \emph{Biometrika}, 63\penalty0 (3):\penalty0 581--592, 1976.

\bibitem[Saria(2018)]{saria2018individualized}
Suchi Saria.
\newblock Individualized sepsis treatment using reinforcement learning.
\newblock \emph{Nature medicine}, 24\penalty0 (11):\penalty0 1641, 2018.

\bibitem[Sharafoddini et~al.(2019)Sharafoddini, Dubin, Maslove, and
  Lee]{sharafoddini2019new}
Anis Sharafoddini, Joel~A Dubin, David~M Maslove, and Joon Lee.
\newblock A new insight into missing data in intensive care unit patient
  profiles: Observational study.
\newblock \emph{JMIR medical informatics}, 7\penalty0 (1):\penalty0 e11605,
  2019.

\bibitem[Sutton and Barto(2018)]{sutton2018reinforcement}
Richard~S Sutton and Andrew~G Barto.
\newblock \emph{Reinforcement learning: An introduction}.
\newblock MIT press, 2018.

\bibitem[Thomas et~al.(2015)Thomas, Theocharous, and
  Ghavamzadeh]{thomas2015high}
Philip~S Thomas, Georgios Theocharous, and Mohammad Ghavamzadeh.
\newblock High-confidence off-policy evaluation.
\newblock In \emph{Twenty-Ninth AAAI Conference on Artificial Intelligence},
  2015.

\bibitem[Vandiver and Vondracek(2012)]{vandiver2012antifactor}
Jeremy~W Vandiver and Thomas~G Vondracek.
\newblock Antifactor xa levels versus activated partial thromboplastin time for
  monitoring unfractionated heparin.
\newblock \emph{Pharmacotherapy: The Journal of Human Pharmacology and Drug
  Therapy}, 32\penalty0 (6):\penalty0 546--558, 2012.

\bibitem[Voloshin et~al.(2019)Voloshin, Le, and Yue]{Voloshin2019}
Cameron Voloshin, Hoang~M Le, and Yisong Yue.
\newblock Empirical analysis of off-policy policy evaluation for reinforcement
  learning.
\newblock \emph{Real-world Sequential Decision Making workshop at ICML 2019.},
  2019.
\newblock URL \url{https://realworld-sdm.github.io/paper/34.pdf}.

\bibitem[Yu et~al.(2019)Yu, Liu, and Nemati]{yu2019reinforcement}
Chao Yu, Jiming Liu, and Shamim Nemati.
\newblock Reinforcement learning in healthcare: A survey.
\newblock \emph{arXiv preprint arXiv:1908.08796}, 2019.

\end{thebibliography}

\documentclass{article}

\PassOptionsToPackage{sort&compress, numbers}{natbib}
     \usepackage[final]{neurips_2019}
    
\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage{hyperref}       
\usepackage{url}            
\usepackage{booktabs}       
\usepackage{amsfonts}       
\usepackage{nicefrac}       
\usepackage{microtype}      
\usepackage{amsmath}

\title{Missingness as Stability: Understanding the Structure of Missingness in Longitudinal EHR data and its Impact on Reinforcement Learning in Healthcare}

\author{%
  Scott L. Fleming$^{1}$, Kuhan Jeyapragasan$^2$, Tony Duan$^2$, Daisy Ding$^1$, \\ \textbf{Saurabh Gombar$^1$, Nigam Shah$^1$, Emma Brunskill$^2$} \\
  $^1$Stanford School of Medicine\\
  $^2$Department of Computer Science, Stanford University\\
  \texttt{scottyf@stanford.edu} \\
}

\begin{document}

\maketitle

\begin{abstract}
There is an emerging trend in the reinforcement learning for healthcare literature. In order to prepare longitudinal, irregularly sampled, clinical datasets for reinforcement learning algorithms, many researchers will resample the time series data to short, regular intervals and use last-observation-carried-forward (LOCF) imputation to fill in these gaps. Typically, they will not maintain any explicit information about which values were imputed. In this work, we (1) call attention to this practice and discuss its potential implications; (2) propose an alternative representation of the patient state that addresses some of these issues; 
and (3) demonstrate in a novel but representative clinical dataset that our alternative representation yields consistently better results for achieving optimal control, as measured by off-policy policy evaluation, compared to representations that do not incorporate missingness information.
\end{abstract}

\section{Introduction}

Recent work in reinforcement learning (RL) for health-related problems has highlighted some of the promises and pitfalls of the RL approach for clinical data \cite{komorowski2018artificial, jeter2019does, saria2018individualized, peng2018improving, gottesman2018evaluating, yu2019reinforcement}. An important but perhaps understated fact throughout this literature is that how one models the patient health state matters: policies learned under one patient state representation may differ dramatically from another, even though the two policies are ultimately learned from the same underlying raw data \cite{gottesman2019guidelines}. An added difficulty is that traditional methods in reinforcement learning are not well-adapted to missingness and the irregular sampling frequencies so commonly observed in clinical data \cite{pivovarov2014identifying, sutton2018reinforcement}. 
RL methods offer a promising framework for optimizing sequences of treatment decisions in the clinic, but these methods will only succeed if they have access to adequate representations of the patient state.

One emerging trend in the recent offline reinforcement learning for healthcare literature is to preprocess the clinical data by discretizing and resampling observations into regular time intervals using ``Last-Observation-Carried-Forward'' (LOCF) imputation \cite{data2016secondary}. In this regime, one substitutes a missing value (e.g. heart rate) with the last observed value for that variable \cite{gelman2006data}. However, few of the reinforcement learning for healthcare papers which utilize LOCF imputation 
mention any use of 
methods 
to leverage missingness information \cite{Jagannatha2018, lopez2019deep, nemati2016optimal, prasad2017reinforcement}. Here, we explore the impact of this choice and potential alternatives. Our results suggest that (1) dynamics of the patient health state differ significantly between periods of more frequent missingness and less frequent missingness, and (2) incorporating missingness information into the patient state representation can improve policy performance, as measured by off-policy policy evaluation (OPPE). 

\section{Background and Related Work}

While researchers have rigorously analyzed the theory of missing data and its implications \cite{rubin1976inference, roderick2002statistical}, missingness in electronic health record (EHR) data specifically is less well studied. 
Early indications suggest, however, that missingness information is a critical component of any adequate patient time series representation. 
This may be because missingness can provide insight into the true (but unobserved) value of the missing entity itself as well as more general information about the patient's health state. Supporting this idea, \cite{beaulieu2018characterizing} explored patterns of missingness in an EHR dataset and found that a substantial portion of common laboratory values appeared \emph{not} to be Missing Completely at Random (where 
observed/missing data would have the same distribution), but rather Missing at Random (where systematic differences between the missing and observed values can occur when considered in isolation, but these differences are fully explained by other observed variables). 
This has important implications. 
\cite{che2018recurrent} found that rates of missingness for many variables in the MIMIC-III dataset were highly correlated with ICD-9 and diagnosis categories. 
\cite{afessa2005influence} found that patients with higher rates of missingness in variables related to an Acute Physiology Score (APS) experienced lower mortality rates compared to those with lower rates of missingness, even after controlling for patients' severity of illness. 

Why might missing values appear in a patient time series extracted from the EHR, and how might they give rise to the above phenomena? We suggest at least two potential mechanisms for missingness, each informative in its own way. The first essentially arises as an artifact of binning a multidimensional time series into discrete windows: if one laboratory test is administered every 24 hours, another every 6 hours, and we bin the patient time series into 4-hour windows, then inevitably some windows will have missing entries for the two tests. Note, however, that if a laboratory test is administered more frequently in sick individuals than in healthy individuals, then rates of missingness for the test will be higher in healthy (less frequently tested) individuals under a fixed binning scheme. Missingness in this context may therefore be a proxy for testing frequency, which in turn can be informative of a patient's overall health. An illustrative example where more frequent sampling (and, consequently, lower rates of missingness in the medical record) is associated with poorer health is given by \cite{duff2018frequency}. 

A second mechanism of missingness arises when a particular laboratory test is only administered (or not administered) if the ordering physician has a suspicion about some aspect of the patient's health. For example, ICU doctors will often order an erythrocyte sedimentation rate (ESR) test on a patient for whom they suspect inflammation or infection \cite{fincher1986clinical, brigden1999clinical}. The \emph{absence} of an ESR value in the EHR, then, could indicate that the clinician did \emph{not} suspect infection or inflammation. This gives additional information about the patient's health state. Illustrating how broad and systematic this mechanism of missignness can be in the EHR, \cite{agniel2018biases} found that just the presence of a laboratory test in a patient's EHR - completely ignoring its actual value - was significantly associated with odds of survival for 86\% of the 272 common clinical laboratory tests they analyzed.

In aggregate, 
current literature supports the notion of "informative missingness" \cite{che2018recurrent, rubin1976inference}, i.e. that missingness itself carries important information about both the value of the missing entity and the patient state. Given the recent attention on reinforcement learning methods as an approach for analyzing longitudinal EHR data, an important lingering question concerns how exactly one should handle missing values in these types of studies. Several approaches from the non-RL literature deserve consideration. \cite{che2018recurrent} developed a unique LSTM structure that incorporates information about a variable's missingness and time since last observation and demonstrated that their model improves predictive performance. \cite{agor2019value},  \cite{sharafoddini2019new}, and \cite{lipton2016directly} all demonstrate that even just the simple practice of including in the feature set an indicator to identify when imputed variables were originally missing can significantly improve performance of predictive models for critical patient outcomes. In this paper, for the sake of illustration and to encourage adoption in practice, we lean toward simple methods over relatively complex ones for incorporating missingness information into the patient state representation and therefore take the latter approach. To the best of our knowledge, ours is the first work to evaluate the impact of including a post-imputation indicator for missingness in the representation of the patient state for applications of reinforcement learning 
to healthcare data.

\section{Dataset}

We extracted EHR data from the STRIDE database \cite{lowe2009stride} for 8,983 ICU patients who underwent unfractionated heparin (UFH) anticoagulation therapy between 2012 and 2018. UFH is an important anticoagulant delivered to ICU patients who are at risk of developing venous thromboembolism (VTE) -- blood clots that can form in a patient’s veins and block the flow of blood to critical organs. In consultation with a clinical care team, all laboratory values deemed relevant to optimal UFH dosing decisions were included in our analysis. Observations in the dataset were defined based on the times at which a physician simultaneously requested an anti-Xa chromogenic assay (anti-Xa) and activated partial thromboplastin time (aPTT) laboratory tests, as these two tests are the standard protocol for UFH monitoring at Stanford Hospital. These orders are frequently made concurrently with a suite of other laboratory tests relevant to the task of UFH dosing, including those listed in Table \ref{mann_whitney}.

The example of UFH dosing is representative of a broader class of optimal control problems often encountered in clinical settings. If a doctor gives a patient too much UFH, they can put the patient at risk of excessive and dangerous bleeding. If the doctor does not give the patient enough UFH, they may increase the patient’s risk of developing a VTE \cite{hylek2003challenges}. For this reason, doctors typically try to titrate a patient’s UFH dosage so as to 
maintain the values of laboratory tests which measure clotting tendency, namely aPTT and anti-Xa, within target therapeutic ranges \cite{baglin2006guidelines}. Abnormalities in liver function, however, as measured by tests like ALT, AST, and TBILI (see Table \ref{mann_whitney}) may alter metabolism of UFH. This can lead to deceptive or unstable aPTT values \cite{vandiver2012antifactor} as well as other more serious complications \cite{price2013discordant}. For this reason, the presence of a lab value like TBILI in a patient's EHR may suggest that the patient is unstable or that the doctor is concerned about more serious underlying physiological issues. Vice versa, the absence of such a lab value may indicate relative stability.

\begin{table}
  \caption{Mann-Whitney U Test evaluating whether aPTT dynamics are significantly different under concurrent laboratory test missingness vs. non-missingness.  (\textbf{bold}: significant after family-wise error rate control 
  ). Only significant results are shown here; see Table \ref{mann_whitney_appendix} in Appendix for full results.}
  \label{mann_whitney}
  \centering
  \begin{tabular}{llll}
    \toprule
    \cmidrule(r){2-3}
    Concurrent Lab, $L$ & Median $X_{L, missing}$ & Median $X_{L, not \ missing}$ & $p$-value \\
    \midrule
    Prothrombin Time (PT)      & 27.10 & 35.44 & \textbf{1.61e-10} \\
    International Normalized Ratio (INR)     & 27.10 & 35.44 & \textbf{1.61e-10} \\
    Total Bilirubin (TBILI)   & 32.71 & 36.21 & \textbf{2.01e-4} \\
    Aspartate Aminotransferase (AST)     & 32.67 & 36.24 & \textbf{2.12e-4} \\
    Alanine Aminotransferase (ALT)     & 32.71 & 36.23 & \textbf{2.57e-4} \\
    Erythrocyte Sedimentation Rate (ESR)     & 34.99 & 39.13 & \textbf{2.20e-5} \\
    \bottomrule
  \end{tabular}
\end{table}

\section{Models and Methods}

\paragraph{Statistical testing.} Given that aPTT values in particular are known to be susceptible to aberrations in liver function, we first tested the hypothesis of whether the dynamics of aPTT, one of the lab values that clinicians would like to keep in control, are significantly different when one of the concurrently-measured laboratory tests is missing vs. not missing. In order to do this, we computed:
\begin{align}
    X_{L, missing} = \{| \textnormal{aPTT}_{t + 1} - \textnormal{aPTT}_{t - 1}| : L_{t}\ \mathrm{missing}\} \\
    X_{L, not\ missing} = \{| \textnormal{aPTT}_{t + 1} - \textnormal{aPTT}_{t - 1}| : \ L_{t}\ \mathrm{not}\ \mathrm{missing}\}
\end{align}

where $L_t$ represents the value of laboratory test $L$ (e.g. total bilirubin) corresponding to an observation taken at time $t$, $\textnormal{aPTT}_{t-1}$ represents the observed aPTT value one observation before this time point, and $\textnormal{aPTT}_{t + 1}$ represents the observed aPTT value one observation after. We then performed a one-sided Mann-Whitney U test \cite{mann1947test}, testing against the null hypothesis that a randomly drawn observation from $X_{L, missing}$ is equally likely to be less than or greater than a randomly drawn observation from $X_{L, not\ missing}$. 
Significant findings from this analysis (after controlling for multiple hypothesis tests with Bonferroni correction 
) can be found in Table \ref{mann_whitney}, with full results in the Appendix (Table \ref{mann_whitney_appendix}).


\paragraph{Reinforcement learning.}

We assessed the impact of including missingness information in the patient state representation on the performance of both Off-Policy Policy Optimization (OPPO) with Fitted Q-Iteration and Off-Policy Policy Evaluation (OPPE) with several Inverse Propensity Scoring (IPS) methods. We defined a reward at each step of the trajectory based on whether or not a patient's aPTT and anti-Xa values fell within what is considered to be a ``therapeutic range''. Both aPTT and anti-Xa measure the patient's blood's propensity to coagulate inappropriately (e.g. too slowly or too quickly). Clinical guidelines suggest that therapeutic ranges are 40 to 80 seconds for aPTT and 0.3 to 0.7 IU/mL for anti-Xa. In each step of a trajectory we define the reward as a sum of -1 for every observation irrespective of the control region (to penalize staying in the ICU), -1 if the patient was not in the therapeutic range for anti-Xa, and -1 if the patient was not in the therapeutic range for aPTT. 

We split our dataset into 5 folds (randomized by patient ID so that all of a patient's visits in the clinic were in the same fold) and performed cross-validation on our entire imputation, OPPO, and OPPE pipeline for each fold \cite{fushiki2011estimation, hastie2005elements}. Thus, for each of the 5 splits of the dataset into training and validation folds, we performed the following steps: (1) for every variable that exhibited missingness in the raw dataset, we added a binary indicator to the patient state representation whose value took on 1 if the corresponding variable was originally missing and 0 otherwise; (2) we performed LOCF imputation on all missing values in the dataset for which a prior value in the patient's trajectory was available; (3) for missing values with no prior value available, we used iterative multivariate imputation, learning the imputation model from the training folds and applying that model to the heldout validation fold \cite{buuren2010mice, buck1960method, scikit-learn}; (4) we learned an optimal policy from the training folds using Fitted Q-Iteration with an Extra-Trees regressor \cite{ernst2005tree}; (5) we evaluated that policy on the held-out validation fold using standard importance sampling (IS), standard weighted importance sampling (WIS), step-wise IS, and step-wise WIS  methods \cite{jiang2015doubly}. 
We used these Inverse Propensity Scoring (IPS) methods for OPPE because they have theoretical guarantees of consistency under assumptions of coverage and nonconfounding and are popular in OPPE literature \cite{Voloshin2019}. 
For each split of the dataset, we performed OPPO with and without the missingness information indicator followed by OPPE with and without the missingness indicator. We do not, however, report evaluation metrics for OPPO with a missingness indicator followed by OPPE without a missingness indicator because of potential confounding \cite{miettinen1974confounding}.

\begin{table}
  \caption{Impact of Including Missingness Information in Patient State Representation on Off-Policy Policy Optimization (OPPO) and Off-Policy Policy Evaluation (OPPE). Numbers reported are the average $\pm$ 1 standard deviation of OPPE estimates over the validation folds from 5-fold CV.}
  \label{ope_heparin}
  \centering
  \begin{tabular}{cccccc}
    \toprule
    \multicolumn{2}{c}{Missingness Information Used?} & \multicolumn{4}{c}{Expected Return (Off-Policy Evaluation)}                \\
    \cmidrule(r){1-2}\cmidrule(r){3-6} 
    OPPO & OPPE & Standard IS & Standard WIS & Stepwise IS & Stepwise WIS\\
    \cmidrule(r){1-2}\cmidrule(r){3-6} 
    No & No  & -0.547 $\pm$ 0.074 & -2.159 $\pm$ 0.236 & -1.137 $\pm$ 0.097 & -3.843 $\pm$ 0.156 \\
    
    No & Yes & -0.538 $\pm$ 0.076  & -2.155 $\pm$ 0.246 & -1.118 $\pm$ 0.100 & -3.838 $\pm$ 0.178\\
    
    
    Yes & Yes & \textbf{-0.504 $\pm$ 0.086}  & \textbf{-2.049 $\pm$ 0.253} & \textbf{-1.046 $\pm$ 0.108} & \textbf{-3.710 $\pm$ 0.232} \\
    \bottomrule
  \end{tabular}
\end{table}

\section{Results and Discussion}

Table \ref{mann_whitney} shows the results of our test for whether the dynamics of aPTT differs under missingness of concurrent laboratory values. 
For six laboratory tests we found that the absolute difference between aPTT values measured just before and just after an aPTT/anti-Xa observation was significantly less, on average, when the the concurrent lab test was missing compared to when it was not missing. Notably, half of the concurrent labs found to have this significant effect (namely TBILI, AST, ALT) are typically used to asses liver damage \cite{newsome2018guidelines}, which is known to disrupt normal aPTT dynamics \cite{ha2016anticoagulation}. These findings suggest that missingness may be a proxy for aPTT test stability over time, and could thus be important when learning to make optimal treatment decisions based on patient state dynamics. 

Table \ref{ope_heparin} shows the effect of including missingness information into the patient state space for OPPO and OPPE. As a comparison, the average discounted rewards under the clinician's data-generating policy was -7.656 $\pm$ 0.268. 
On average, incorporating missingness into the patient state space for both OPPO and OPPE improved the estimated reward under every IPS method analyzed.
The effects appear to be additive in the sense that incorporating missingness into the state representation for OPPE improved estimated performance, and additionally incorporating missingness into 
OPPO increased estimated performance even further. Although these findings were provocatively consistent across IPS methods, they were not significant, likely due to known issues of high variance in OPPE estimators \cite{thomas2015high}. In accordance with our discussion in Section 2 about "informative missingness", our results suggest that missingness in our dataset may carry important information about the patient's health state, and that leveraging this missingness information can potentially improve estimated reward under OPPO and OPPE. Researchers applying reinforcement learning to healthcare datasets may benefit from this simple augmentation to the patient state representation. The code used to run our analyses can be found at \url{github.com/scottfleming/rl-missingness}. 

\medskip

\bibliography{references}
\pagebreak
\section{Appendix}
\begin{table}[h]
  \caption{Mann-Whitney U Test evaluating whether aPTT dynamics are significantly different under concurrent laboratory test missingness vs. non-missingness.  (\textbf{bold}: significant after Bonferroni correction for controlling the family-wise error rate 
  )}
  \label{mann_whitney_appendix}
  \centering
  \begin{tabular}{llll}
    \toprule
    \cmidrule(r){2-3}
    Concurrent Lab, $L$ & Median $X_{L, missing}$ & Median $X_{L, not \ missing}$ & $p$-value \\
    \midrule
    Prothrombin Time (PT)      & 27.10 & 35.44 & \textbf{1.61e-10} \\
    International Normalized Ratio (INR)     & 27.10 & 35.44 & \textbf{1.61e-10} \\
    Total Bilirubin (TBILI)   & 32.71 & 36.21 & \textbf{2.01e-4} \\
    Creatinine (CR)      & 27.49 & 35.27 & 2.22e-2 \\
    Aspartate Aminotransferase (AST)     & 32.67 & 36.24 & \textbf{2.12e-4} \\
    Alanine Aminotransferase (ALT)     & 32.71 & 36.23 & \textbf{2.57e-4} \\
    Platelet Count (PLT)     & 42.48 & 35.18 & 0.988 \\
    C-reactive Protein (CRP)     & 35.15 & 35.53 & 0.965 \\
    Red Cell Distribution Width (RDW)     & 33.67 & 35.19 & 0.561 \\
    Hemoglobin (HGB)     & 40.14 & 35.18 & 0.372 \\
    White Blood Cell Count (WBC)     & 33.67 & 35.19 & 0.561 \\
    Fibrinogen (FBN)     & 34.27 & 39.84 & 0.324 \\
    Erythrocyte Sedimentation Rate (ESR)     & 34.99 & 39.13 & \textbf{2.20e-5} \\
    \bottomrule
  \end{tabular}
\end{table}

\end{document}